OpenStudy (chris1234):

Distributive Property: -6(t-1)

hartnn (hartnn):

a(b+c) =ab+ac so, -6(t-1) = -6*t + (-6)(-1) now simplify

OpenStudy (anonymous):

-6t+6

OpenStudy (anonymous):

-6t+6? what do we have to do

OpenStudy (anonymous):

-6t+6 is the answer

OpenStudy (anonymous):

You just multiply -6 times t and you get -6t. Then you multiply -6 times -1 you get positive 6 So when you put that together you get -6t+6

OpenStudy (anonymous):

yeah i know that now but thanks